OpenStudy (anonymous):

The sum of two consecutive odd integers is 68. hat is the smaller number?

ganeshie8 (ganeshie8):

try 31+33 try 33+35 try 35+37

OpenStudy (kc_kennylau):

Let the smaller integer be n, and the larger integer be n+2.

OpenStudy (kc_kennylau):

How to express your question in equation?

ganeshie8 (ganeshie8):

or express it as equation and do it in royal way ;)

OpenStudy (anonymous):

here are the answers i have to choose from A.29 B. 31 C. 33 D.35

OpenStudy (wolf1728):

An odd integer is expressed as 2x + 1 So the 2 consecutive odd integers are 2x +1 and 2x +3 2x +1 + 2x + 3 = 68 4x + 4 = 68 4x = 64 x = 16 2x +1 = 33 2 x + 3 = 35 Smaller number is 33

OpenStudy (anonymous):

(2x+1)+(2x+3)=68 4x+4=68 4x=64 x=16 therefore smaller number= 2x+1 =2(16)+1 =33

OpenStudy (anonymous):

lol did working then u posted urs first

OpenStudy (kc_kennylau):

Let the smaller integer be n, and the larger integer be n+2. n+(n+2)=68 2n+2=68 2n=66 n=33 Therefore the smaller integer is 33.

OpenStudy (kc_kennylau):

So much simpler :P

OpenStudy (wolf1728):

well great minds think alike !!

OpenStudy (kc_kennylau):

nope my mind ain't great great is your mind :)

OpenStudy (wolf1728):

Sounds a bit like Yoda

OpenStudy (kc_kennylau):

yoda goes like this "great, your mind is" xD